OpenStudy (kirstentaylort):

The sales tax in Bill's state is 6%. Bill bought a Scion having a sales tax of $820. What was the cost of the car? Round to the nearest dollar. A. $13,665 B. $13,666 C. $13,668 D. $13,667

OpenStudy (mathstudent55):

Are you familiar with proportions?

OpenStudy (kirstentaylort):

not at all

OpenStudy (kirstentaylort):

maybe but i would have to refresh my memory

OpenStudy (mathstudent55):

Ok. Let's use percents then. The sales tax is 6%. The car before tax cost an unknown amount which we will call x. The problem is telling us that 6% of x is $820 Ok so far?

OpenStudy (kirstentaylort):

okay

OpenStudy (mathstudent55):

We translate the last statement into an equation and we solve the equation. 6% of x = $820 In math, "of" usually means multiplication. Also, to change a percent to a decimal, divide the percent by 100. 6% = 0.06 6% of x = 820 6% * x = 820 0.06x = 820 Do you understand the steps above?

OpenStudy (mathstudent55):

Now we need to solve for x. Since x is being multiplied by 0.06, and we want x alone, we divide both sides by 0.06. \(\dfrac{0.06x}{0.06} = \dfrac{820}{0.06} \) \(x = \dfrac{820}{0.06} \)